# Qudit-native measurement protocol for dynamical correlations using Hadamard tests

### v1

@@ Appendix B Variance of the Hadamard-test
Finally, identifying the expectation value $\langle\hat{\Pi}_{0}\rangle$ with the probability $P_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})$ to measure $|0\rangle$ on the ancilla, concludes the proof.
@@
In order to arrive from Eq. 7 at the a priori upper bound $\operatorname{Var}[C^{\pm}_{AB}(t_{1},t_{2})]\leq 4\lVert\hat{A}\rVert^{2}\lVert\hat{B}\rVert^{2}$, we have used the fact that the function $f(x)=x(1-x)$ has a maximum value $f(x=1/2)=1/4$ in the interval $[0,1]$.
@@ Appendix C Decomposition for many-body observables
where each $\hat{S}^{\alpha_{k}}_{k}$ with $\alpha_{k}\in\{x,y,z\}$ is a spin operator. The unitary $\hat{W}_{X}$ can then be calculated using the unitary decomposition of each of the (one-body) spin operators:
@@
$\hat{W}_{X}=\hskip 10.00002pt\sum_{\mathclap{\left\{\hat{V}_{k}=\hat{W}_{\alpha_{k}},\hat{W}_{\alpha_{k}}^{\dagger}\right\}}}\hat{V}_{1}\otimes\cdots\otimes\hat{V}_{n}.$ (22)
In this case, the experimental implementation of the controlled unitaries in the Hadamard-test simplifies since we only have to deal with a product of two-body operations. However, the number of Hadamard tests to be performed increases exponentially with the length of the string in Eq. (20).
For a general multi-qudit observable, similar representations as strings of one-qudit observables can be employed to simplify the experimental implementation of controlled operations in the Hadamard-test.

### v2

@@ Appendix B Variance of the Hadamard test
We proceed with the proof of Eq. 7. To this end, we rewrite Eq. 6 by expressing the probability $P_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})$ of measuring $|0\rangle$ on the ancilla as the expectation value of the projector $\hat{\Pi}_{0}=|0\rangle\langle 0|\otimes\mathds{1}$:
$C_{AB}^{\pm}(t_{1},t_{2})=\big{\lVert}\hat{A}\big{\rVert}\big{\lVert}\hat{B}\big{\rVert}\\ \times\sum_{\mathclap{V,V^{\prime}\in\{W,W^{\dagger}\}}}\Big{(}\big{\langle}\Psi_{V_{A}V_{B}^{\prime}}^{\pm}(t_{1},t_{2})\,\big{|}\,\mathopen{}\hat{\Pi}_{0}\,\big{|}\,\mathopen{}\Psi_{V_{A}V_{B}^{\prime}}^{\pm}(t_{1},t_{2})\big{\rangle}-\frac{1}{2}\Big{)}\,.$ (21)
Since the measurements of the four Hadamard-test configurations with controlled unitaries $\hat{V}_{A}\in\{\hat{W}_{A},\hat{W}^{\dagger}_{A}\}$ and $\hat{V}_{B}\in\{\hat{W}_{B},\hat{W}^{\dagger}_{B}\}$ are independent, standard propagation of variance yields [67]
$\operatorname{Var}[C^{\pm}_{AB}(t_{1},t_{2})]=\big{\lVert}\hat{A}\big{\rVert}^{2}\big{\lVert}\hat{B}\big{\rVert}^{2}\sum_{\mathclap{V,V^{\prime}\in\{W,W^{\dagger}\}}}\operatorname{Var}[\hat{\Pi}_{0}]^{\pm}_{V_{A}V_{B}^{\prime}}(t_{1},t_{2})\,.$ (22)
If the Hadamard test with controlled unitaries $\hat{V}_{A}$ and $\hat{V}_{B}$ is repeated $M_{V_{A}V_{B}}$ times, the probability $P_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})$ can be estimated with variance
$\operatorname{Var}[\hat{\Pi}_{0}]^{\pm}_{V_{A}V_{B}}(t_{1},t_{2})\\ \begin{split}=\frac{1}{M_{V_{A}V_{B}}}\Big{[}&\big{\langle}\Psi_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})\,\big{|}\,\mathopen{}\hat{\Pi}_{0}^{2}\,\big{|}\,\mathopen{}\Psi_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})\big{\rangle}\\ &-\big{\langle}\Psi_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})\,\big{|}\,\mathopen{}\hat{\Pi}_{0}\,\big{|}\,\mathopen{}\Psi_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})\big{\rangle}^{2}\Big{]}\,.\end{split}$ (23)
For the projector $\hat{\Pi}_{0}$, the simple identity $\hat{\Pi}_{0}^{2}=\hat{\Pi}_{0}$ holds, and thus for an arbitrary state
$\langle\hat{\Pi}_{0}^{2}\rangle-\langle\hat{\Pi}_{0}\rangle^{2}=\langle\hat{\Pi}_{0}\rangle\big{(}1-\langle\hat{\Pi}_{0}\rangle\big{)}\,.$ (24)
Finally, identifying the expectation value $\langle\hat{\Pi}_{0}\rangle$ with the probability $P_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})$ concludes the proof.
@@
In order to arrive from Eq. 7 at the a priori upper bound $\operatorname{Var}[C^{\pm}_{AB}(t_{1},t_{2})]\leq M^{-1}\lVert\hat{A}\rVert^{2}\lVert\hat{B}\rVert^{2}$, we used the fact that the function $f(x)=x(1-x)$ has a maximum value $f(x=1/2)=1/4$ in the interval $[0,1]$.
@@ Appendix C Decomposition for many-body observables
where each $\hat{S}^{\alpha_{k}}_{k}$, with $\alpha_{k}\in\{x,y,z\}$, is a spin operator. The unitary $\hat{W}_{X}$ can then be calculated using the unitary decomposition of each of the (one-body) spin operators:
@@
$\hat{W}_{X}=\hskip 10.00002pt\sum_{\mathclap{\left\{V_{k}=W_{\alpha_{k}},W_{\alpha_{k}}^{\dagger}\right\}}}\hat{V}_{1}\otimes\cdots\otimes\hat{V}_{n}\,.$ (27)
In this case, the experimental implementation of the controlled unitaries in the Hadamard-test simplifies since we have to deal with only a product of two-body operations. However, the number of Hadamard tests to be performed increases exponentially with the length of the string in Eq. (25).
For a general multi-qudit observable, representations similar to strings of one-qudit observables can be employed to simplify the experimental implementation of controlled operations in the Hadamard test. In practice, however, an approximate decomposition such as the one resulting from the LCU procedure [54] may be more efficient in terms of circuit depth for implementations on NISQ devices.